3- Leaf Company is about to issue a bond with semiannual coupon payments, an annual coupon rate of 11%, and par value of $1,000. The yield to maturity for this bond is 9%.

What is the price of the bond if it matures in 30 years?
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

M = $1000, n = 30 * 2 = 60 semi-annual periods, i = 9%/2 = 4.5% (semi-annually), C = 11% * $1000/2 = $55

1- P = 55+ -- 1:05 1000 P = 55* (1+0.045)60 0.045 1000 (1 +0.045)60

P = $1,135.09 + $71.29

P = $1,206.38